Inactive Tutor answered 01/08/20
Let us start by considering the probability of each occurrence separately (treat as independent events, since the previous outcomes do not affect later ones):
Head on the First flip: P = 1/2
Tail on the Second flip: P = 1/2
Head on the Third flip: P = 1/2
With these independent events, the probability of all events happening is simply the product: 1/2 * 1/2*1/2 = 1/8